# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ B.2 Extension to the General Risk Loss $\ell$ under the Nonparametric Setting
Example 3 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{G}\asymp\mathcal{F}$ FAIR-Linear Theorem B.10
Example 7 $\mathcal{H}_{\mathtt{nn}}(d,L_{g},N_{g},B_{g})$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\asymp\mathcal{F}$ FAIR-NN Theorem 4.1
Example 4 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{alin}}(d,\phi)$ $\mathcal{G}\ll\mathcal{F}$ FAIR-AugLinear Theorem B.11
Example 5 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\ll\mathcal{F}$ FAIR-NNLinear Theorem B.12
Example 6 $\mathcal{H}_{\mathtt{ann}}(d,L_{g},N_{g},B_{g})$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\ll\mathcal{F}$ FAIR-ANN Theorem B.9
@@ B.5 FAIR-ANN: Bridging Invariance and Additional Structural Knowledge
Condition B.5 (Invariance and Nondegenerate Covariate for FAIR-ANN).
@@
Condition B.6.
@@
Condition B.7 (Identification for FAIR-ANN).
@@
Theorem B.9 (Optimal Rate for FAIR-ANN Least Squares Estimator).
Assume 4.1,4.3, 4.4, and B.5B.7 hold. Assume further that all the conditional moments $\{m^{(e,S)}\}_{e\in\mathcal{E},S\subseteq[d]}$ are $(\beta^{\prime},C^{\prime})$-smooth for some $\beta^{\prime}>0$ and $C^{\prime}>0$, and $\delta_{\mathtt{opt}}=o(1)$. Consider the FAIR-ANN estimator that solves (3.8) with $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ using $\gamma\geq 8\gamma^{\star}_{\mathtt{AN}}$ with
@@
The choice of $N,L$, and the convergence rate align with FAIR-NN with $\alpha^{\star}=\beta^{\star}$. Given the strong structural prior knowledge that the true regression function is additive, FAIR-ANN requires weaker identification condition B.7 and also smaller critical threshold of $\gamma$. In particular, B.7 requires that for any $S$ such that regressing $Y$ on $X_{S\cup S^{\star}}$ via additive models yields biased estimation, there should be either (1) a shift in conditional moments $m^{(e,S)}$ across different environments, or (2) one of the conditional moments $m^{(e,S)}$ is non-additive. This characteristic is called the “double identifiable” property since meeting either of these conditions can consistently estimate $m^{\star}$. Notably, the critical threshold $\gamma^{\star}_{\mathtt{AN}}$ can be smaller than that of the FAIR-NN estimator. A small $\gamma$ can be adopted if either the signal of violating the additive structure or the signal of heterogeneity is strong.
@@ B.6 Theoretical Analysis for Linear $\mathcal{G}$
In this section, we apply our result in Theorem B.2 to the cases where the target regression function $g^{\star}$ is linear. As such, we use linear function class $\mathcal{H}_{\mathtt{lin}}(d)$ as our predictor function class $\mathcal{G}$. Our theorem suggests that enhancing the potential approximation ability of the discriminator function class $\mathcal{F}$ will result in (1) a stronger condition on invariance, and (2) a weaker identification condition and a reduced choice of critical threshold $\gamma^{\star}$.
@@ B.6.1 Linear $\mathcal{F}$
    The data satisfies 4.1 with $|\mathcal{E}|\leq n^{C_{1}}$ for some constant $C_{1}$.
@@
    4.3 holds.
Under B.8 that the covariance matrices are all positive definite, we can define
@@
Condition B.9 (Invariance in Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
@@
Condition B.10 (Identification for Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
@@
Theorem B.10 (Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
Suppose B.8B.10 hold, and we choose
@@
We present the results using truncated function classes, and there exist poly-$\log n$ factors in the non-asymptotic $L_{2}$ error bounds. These are for technical convenience such that we can directly apply our result Theorem B.2 which focuses on uniformly bounded function classes. Indeed, one can use a finer analysis and obtain the $\ell_{2}$ error bound
@@
The obtained results in Theorem B.10 align with (up to $\log(n)$ factors) and offer significant enhancements over Theorem 2 & 3 from Fan et al., (2023). Firstly, the “invariance” condition gets relaxed, we only assume that the noise $\varepsilon^{(e)}$ and the true important variables $X_{S^{\star}}^{(e)}$ are uncorrelated rather than conditional independent across different environments. Meanwhile, the identification condition B.10 exactly matches that in Fan et al., (2023) (refer to Condition 5 therein), and the choice of critical threshold $\gamma^{\star}$ gets reduced as indicated by the inequality in (B.14) and given that $\kappa_{L}=O(1)$. Such an improvement can be attributed to the term $-\frac{1}{2}\{f^{(e)}\}^{2}$ in our minimax regularization that stabilizes the objective. To see this, consider $\beta$ with $\mathrm{supp}(\beta)=S^{\star}$, the population-level EILLS objective can be written as
@@ B.6.2 Augmented Linear $\mathcal{F}$
Under B.11 such that the covariance matrix for $\widetilde{X}$ are positive definite, we can define
@@
Condition B.12 (Invariance in Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
@@
Condition B.13 (Identification for Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
@@
Theorem B.11 (Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
Suppose B.8, B.11B.13 hold, and we choose
@@
    When there are multiple environments $|\mathcal{E}|>1$, the identification condition B.13 is weaker to both the EILLS and FGMM estimators. In particular, a consistent estimate $\beta^{\star}$ is attainable if incorporating variables $x_{j}$ with $\sum_{e\in\mathcal{E}}\mathbb{E}[X_{j}^{(e)}\varepsilon^{(e)}]\neq 0$ will result in either (1) a shift in the best linear predictor across environments or (2) the fitted residuals is strongly correlated with some nonlinear basis. We refer to this property as “double identifiable” property, given satisfying either condition can lead to the consistent estimation of the true parameter. Furthermore, the critical threshold $\gamma^{\star}$ can be smaller than that of the EILLS estimator according to the inequality $\bar{\mathsf{d}}_{\mathtt{LA}}(S)\geq\bar{\mathsf{d}}_{\mathtt{LL}}(S)$. This implies that the estimation is sample efficient, which allows for a small $\gamma$, if either the signal of nonlinear basis or the signal of heterogeneity is strong.
@@
At the same time, it should be noted that the above advantages over the EILLS estimator (linear $\mathcal{F}$) are at the cost of imposing stronger invariance condition B.12, which assures that the noise should not only be uncorrelated with $X_{j}^{(e)}$ but also be uncorrelated with $\phi(X_{j}^{(e)})$ for any $j\in S^{\star}$ and $e\in\mathcal{E}$.
@@ B.6.3 Neural Network $\mathcal{F}$
Condition B.15 (Invariance in Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
@@
Condition B.16 (Identification for Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
@@
Theorem B.12 (Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
Suppose B.8, B.14B.16 hold, and we choose the function classes $\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d,C_{2},C_{2}\sqrt{\log n})$ and $\mathcal{H}_{\mathtt{nn}}(d,\log^{d}n,\log^{d}n,C_{2}\sqrt{\log n})$ with some constant $C_{2}\geq(1\lor\sigma_{x}\lor\sigma_{m})\max_{e\in\mathcal{E},S\subseteq[d]}\|\Sigma^{1/2}\beta^{\star}\|_{2}$. Then, there exists some constant $\widetilde{C}$ that only depends on $(C_{1},C_{2},d,\sigma_{m},C_{m},\sigma_{y},C_{y},\sigma_{x},C_{x})$ such that the FAIR estimator using the above function classes and hyper-parameter $\gamma$ satisfying $\gamma\geq 8\gamma^{\star}_{\mathtt{LN}}=8\sup_{S:\mathsf{b}_{\mathtt{LL}}(S)>0}\mathsf{b}_{\mathtt{LL}}(S)/\bar{\mathsf{d}}_{\mathtt{LN}}(S)$, where
@@
The estimator can be viewed as an advanced version of the one using $\mathcal{F}=\mathcal{H}_{\mathtt{alin}}(d,\phi)$. It leverages neural networks to search for appropriate basis function $\phi$ with strong signals. With the proper choice of the neural network hyper-parameters, the estimator still maintains a parametric optimal rate (up to logarithmic factors). Additionally, it requires a weaker identification condition as described by B.16 and reduced critical threshold $\gamma^{\star}$ according to the inequality $\bar{\mathsf{d}}_{\mathtt{LN}}(S)\geq\bar{\mathsf{d}}_{\mathtt{LA}}(S)$ in Theorem B.12.
@@ C.1 Pseudo-code of the Gradient Descent Ascent Algorithm
14:         Sample $\{u_{i,1},u_{i,2}\}_{i=1}^{d}$ from $\mathrm{Gumbel}(0,1)$
15:         Calculate $a=(a_{1},\ldots,a_{d})$ with $a_{j}=V(w_{j},\tau_{t},u_{j,1},u_{j,2})$, where $V(\cdot)$ is defined in (6.1).
@@
22:         Sample $\{u_{i,1},u_{i,2}\}_{i=1}^{d}$ from $\mathrm{Gumbel}(0,1)$
23:         Calculate $a=(a_{1},\ldots,a_{d})$ with $a_{j}=V(w_{j},\tau_{t},u_{j,1},u_{j,2})$, where $V(\cdot)$ is defined in (6.1).
@@ C.2 Detailed Simulation Configuration
C.2.1 Linear Model with $p=15$
@@ C.2.2 Nonlinear Model
For the FAIR-NN implementation using Gumbel approximation, we also run gradient descent ascent using the Adam optimizer using a learning rate of 1e-3, batch size $64$. The number of iterations is $70k$ for $m_{1}^{\star}$ and $80k$ for $m_{2}^{\star}$. In each iteration, one gradient descent update of the neural network parameters in $g$ and the Gumbel logits parameter $w$ is conducted followed by three gradient ascent updates of the neural network parameters in $f^{0}$ and $f^{(1)}$. We also use fixed $\gamma=36$. The implementation details for the estimators are:
@@
For FAIR-GB, we report the estimated MSE for the model in the last iteration. For other estimators, we also run gradient descent using Adam optimizer for 10k iterations. We report the estimated MSE for the model with early stopping regularization: that is, we report the estimated MSE of the model that has the smallest validation error, and the validation data is sampled independently and identically to the training data with sample size $n_{\mathrm{valid}}=\lfloor 3n/7\rfloor$.

### v2

@@ B.4 Applications of Theorem 4 and Connection to the Predecessors
Example 3 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{G}\asymp\mathcal{F}$ FAIR-Linear Theorem 8
Example 7 $\mathcal{H}_{\mathtt{nn}}(d,L_{g},N_{g},B_{g})$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\asymp\mathcal{F}$ FAIR-NN Theorem 1
Example 4 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{alin}}(d,\phi)$ $\mathcal{G}\ll\mathcal{F}$ FAIR-AugLinear Theorem 9
Example 5 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\ll\mathcal{F}$ FAIR-NNLinear Theorem 10
Example 6 $\mathcal{H}_{\mathtt{ann}}(d,L_{g},N_{g},B_{g})$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\ll\mathcal{F}$ FAIR-ANN Theorem 7
@@ B.5 FAIR-ANN: Bridging Invariance and Additional Structural Knowledge
Condition 13 (Invariance and Nondegenerate Covariate for FAIR-ANN).
@@
Condition 14 (Boundedness in Nonparametric Regression).
Condition 15.
@@
Condition 16 (Identification for FAIR-ANN).
@@
Theorem 7 (Optimal Rate for FAIR-ANN Least Squares Estimator).
Assume 7,8, and 1316 hold. Assume further that all the conditional moments $\{m^{(e,S)}\}_{e\in\mathcal{E},S\subseteq[d]}$ are $(\beta^{\prime},C^{\prime})$-smooth for some $\beta^{\prime}>0$ and $C^{\prime}>0$, and $\delta_{\mathtt{opt}}=o(1)$. Consider the FAIR-ANN estimator that solves (B.2) with $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ using $\gamma\geq 8\gamma^{\star}_{\mathtt{AN}}$ with
@@
The choice of $N,L$, and the convergence rate align with FAIR-NN with $\alpha^{\star}=\beta^{\star}$. Given the strong structural prior knowledge that the true regression function is additive, FAIR-ANN requires weaker identification condition 16 and also smaller critical threshold of $\gamma$. In particular, 16 requires that for any $S$ such that regressing $Y$ on $X_{S\cup S^{\star}}$ via additive models yields biased estimation, there should be either (1) a shift in conditional moments $m^{(e,S)}$ across different environments, or (2) one of the conditional moments $m^{(e,S)}$ is non-additive. This characteristic is called the “double identifiable” property since meeting either of these conditions can consistently estimate $m^{\star}$. Notably, the critical threshold $\gamma^{\star}_{\mathtt{AN}}$ can be smaller than that of the FAIR-NN estimator. A small $\gamma$ can be adopted if either the signal of violating the additive structure or the signal of heterogeneity is strong.
@@ B.6 Theoretical Analysis for Linear $\mathcal{G}$
In this section, we apply our result in Theorem 4 to the cases where the target regression function $g^{\star}$ is linear. As such, we use linear function class $\mathcal{H}_{\mathtt{lin}}(d)$ as our predictor function class $\mathcal{G}$. Our theorem suggests that enhancing the potential approximation ability of the discriminator function class $\mathcal{F}$ will result in (1) a stronger condition on invariance, and (2) a weaker identification condition and a reduced choice of critical threshold $\gamma^{\star}$.
@@ B.6.1 Linear $\mathcal{F}$
    The data satisfies 7 with $|\mathcal{E}|\leq n^{C_{1}}$ for some constant $C_{1}$.
@@
    8 holds.
Under 17 that the covariance matrices are all positive definite, we can define
@@
Condition 18 (Invariance in Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
@@
Condition 19 (Identification for Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
@@
Theorem 8 (Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
Suppose 1719 hold, and we choose
@@
We present the results using truncated function classes, and there exist poly-$\log n$ factors in the non-asymptotic $L_{2}$ error bounds. These are for technical convenience such that we can directly apply our result Theorem 4 which focuses on uniformly bounded function classes. Indeed, one can use a finer analysis and obtain the $\ell_{2}$ error bound
@@
The obtained results in Theorem 8 align with (up to $\log(n)$ factors) and offer significant enhancements over Theorem 2 & 3 from Fan et al., (2023). Firstly, the “invariance” condition gets relaxed, we only assume that the noise $\varepsilon^{(e)}$ and the true important variables $X_{S^{\star}}^{(e)}$ are uncorrelated rather than conditional independent across different environments. Meanwhile, the identification condition 19 exactly matches that in Fan et al., (2023) (refer to Condition 5 therein), and the choice of critical threshold $\gamma^{\star}$ gets reduced as indicated by the inequality in (B.15) and given that $\kappa_{L}=O(1)$. Such an improvement can be attributed to the term $-\frac{1}{2}\{f^{(e)}\}^{2}$ in our minimax regularization that stabilizes the objective. To see this, consider $\beta$ with $\mathrm{supp}(\beta)=S^{\star}$, the population-level EILLS objective can be written as
@@ B.6.2 Augmented Linear $\mathcal{F}$
Under 20 such that the covariance matrix for $\widetilde{X}$ are positive definite, we can define
@@
Condition 21 (Invariance in Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
@@
Condition 22 (Identification for Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
@@
Theorem 9 (Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
Suppose 17, 2022 hold, and we choose
@@
    When there are multiple environments $|\mathcal{E}|>1$, the identification condition 22 is weaker to both the EILLS and FGMM estimators. In particular, a consistent estimate $\beta^{\star}$ is attainable if incorporating variables $x_{j}$ with $\sum_{e\in\mathcal{E}}\mathbb{E}[X_{j}^{(e)}\varepsilon^{(e)}]\neq 0$ will result in either (1) a shift in the best linear predictor across environments or (2) the fitted residuals is strongly correlated with some nonlinear basis. We refer to this property as “double identifiable” property, given satisfying either condition can lead to the consistent estimation of the true parameter. Furthermore, the critical threshold $\gamma^{\star}$ can be smaller than that of the EILLS estimator according to the inequality $\bar{\mathsf{d}}_{\mathtt{LA}}(S)\geq\bar{\mathsf{d}}_{\mathtt{LL}}(S)$. This implies that the estimation is sample efficient, which allows for a small $\gamma$, if either the signal of nonlinear basis or the signal of heterogeneity is strong.
@@
At the same time, it should be noted that the above advantages over the EILLS estimator (linear $\mathcal{F}$) are at the cost of imposing stronger invariance condition 21, which assures that the noise should not only be uncorrelated with $X_{j}^{(e)}$ but also be uncorrelated with $\phi(X_{j}^{(e)})$ for any $j\in S^{\star}$ and $e\in\mathcal{E}$.
@@ B.6.3 Neural Network $\mathcal{F}$
Condition 24 (Invariance in Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
@@
Condition 25 (Identification for Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
@@
Theorem 10 (Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
Suppose 17, 2325 hold, and we choose the function classes $\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d,C_{2},C_{2}\sqrt{\log n})$ and $\mathcal{H}_{\mathtt{nn}}(d,\log^{d}n,\log^{d}n,C_{2}\sqrt{\log n})$ with some constant $C_{2}\geq(1\lor\sigma_{x}\lor\sigma_{m})\max_{e\in\mathcal{E},S\subseteq[d]}\|\Sigma^{1/2}\beta^{\star}\|_{2}$. Then, there exists some constant $\widetilde{C}$ that only depends on $(C_{1},C_{2},d,\sigma_{m},C_{m},\sigma_{y},C_{y},\sigma_{x},C_{x})$ such that the FAIR estimator using the above function classes and hyper-parameter $\gamma$ satisfying $\gamma\geq 8\gamma^{\star}_{\mathtt{LN}}=8\sup_{S:\mathsf{b}_{\mathtt{LL}}(S)>0}\mathsf{b}_{\mathtt{LL}}(S)/\bar{\mathsf{d}}_{\mathtt{LN}}(S)$, where
@@
The estimator can be viewed as an advanced version of the one using $\mathcal{F}=\mathcal{H}_{\mathtt{alin}}(d,\phi)$. It leverages neural networks to search for appropriate basis function $\phi$ with strong signals. With the proper choice of the neural network hyper-parameters, the estimator still maintains a parametric optimal rate (up to logarithmic factors). Additionally, it requires a weaker identification condition as described by 25 and reduced critical threshold $\gamma^{\star}$ according to the inequality $\bar{\mathsf{d}}_{\mathtt{LN}}(S)\geq\bar{\mathsf{d}}_{\mathtt{LA}}(S)$ in Theorem 10.
@@ C.1 Pseudo-code of the Gradient Descent Ascent Algorithm
14:         Sample $\{u_{j}\}_{j=1}^{d}=\{(u_{j,1},u_{j,2})\}_{j=1}^{d}$ from $\mathrm{Gumbel}(0,1)$.
15:         Calculate $a=(a_{1},\ldots,a_{d})$ with $a_{j}=V_{\tau_{t}}(w_{j},u_{j})$, where $V(\cdot)$ is defined in (5.2).
@@
22:         Sample $\{u_{j}\}_{j=1}^{d}=\{(u_{j,1},u_{j,2})\}_{j=1}^{d}$ from $\mathrm{Gumbel}(0,1)$.
23:         Calculate $a=(a_{1},\ldots,a_{d})$ with $a_{j}=V_{\tau_{t}}(w_{j},u_{j})$, where $V(\cdot)$ is defined in (5.2).
@@ C.2 Detailed Simulation Configuration
C.2.1 Linear Model with $d=15$
@@ C.2.2 Nonlinear Model
For the FAIR-NN implementation using Gumbel approximation, we also run gradient descent ascent using the Adam optimizer using a learning rate of 1e-3, batch size $64$. The number of iterations is $70k$ for $m_{1}^{\star}$ and $80k$ for $m_{2}^{\star}$. In each iteration, one gradient descent update of the neural network parameters in $g$ and the Gumbel logits parameter $w$ is conducted followed by three gradient ascent updates of the neural network parameters in $f^{(0)}$ and $f^{(1)}$. We also use fixed $\gamma=36$. The implementation details for the estimators are:
@@
For FAIR-GB, we report the estimated MSE for the model in the last iteration. For other estimators, we also run gradient descent using the Adam optimizer for 10k iterations. We report the estimated MSE for the model with early stopping regularization: that is, we report the estimated MSE of the model that has the smallest validation error, and the validation data is sampled independently and identically to the training data with sample size $n_{\mathrm{valid}}=\lfloor 3n/7\rfloor$.
C.3 Details of the Discovery in Real Physical System Application
Data Collection
We directly use the dataset ‘lt_interventions_standard_v1’ released in Gamella et al., (2024).
For the training dataset, given fixed sample size $n$, the data in the first environment $e=0$ is sampled from the experimental setting ‘uniform_reference’. For the second environment $e=1$, a mixture of interventions is applied. To be specific, a weak intervention on the variables $\widetilde{V}_{3},\widetilde{V}_{1},\widetilde{V}_{2},\widetilde{I}_{1},\widetilde{I}_{2}$ with probability $(1/3,1/6,1/6,1/6,1/6)$, respectively. This is equivalent to sample data from the experimental setting ‘t_vis_3_weak’, ‘t_vis_1_weak’, ‘t_vis_2_weak,’, ‘t_ir_1_weak’, ‘t_ir_2_weak’ with weights $(1/3,1/6,1/6,1/6,1/6)$.
For the test data used for evaluation in Fig. 6 (b)–(c), we use the data from the experimental setting ‘t_vis_3_strong’, ‘t_vis_1_strong’, ‘t_vis_2_strong’, ‘t_ir_1_strong’, ‘t_ir_2_strong’. Since there is an out-of-support issue for the intervention, i.e.,
$|\mathrm{Mean}_{\mu_{X,i}}(X)-\mathrm{Mean}_{\bar{\mu}_{n}}(X)|>1.6\cdot\mathrm{Std}_{\bar{\mu}_{n}}(X)$
where $\mu_{X,i}$ is the empirical distribution of $X$ in the experimental setting where strong intervention is intervened on $X$, and $\bar{\mu}_{n}$ is the empirical distribution of $X$ in the training dataset. Thus, we recenter the variable $X$ in the corresponding test intervention environment such that it has the same empirical mean as that in the training dataset.
Explanation on the Equivalent Graph
We regress $\widetilde{I}_{3}$ on $(R,G,B,\theta_{1},\theta_{2},\widetilde{V}_{1},\widetilde{V}_{2},\widetilde{V}_{3},\widetilde{I}_{1},\widetilde{I}_{2})$. There are several hidden confounders, hence there should be an arrow from $\widetilde{V}_{3}$ to $\widetilde{I}_{3}$ and an arrow from $\widetilde{I}_{3}$ to $\widetilde{V}_{3}$ if $\widetilde{V}_{3}$ is not intervened given the existence of hidden confounders $(L_{3,1},L_{3,2})$. Introducing the variable $\widetilde{V}_{3}$ in predicting $\widetilde{I}_{3}$ can increase the predictive power given it can provide additional information of $(L_{3,1},L_{3,2})$. The (equivalent) arrow from $\widetilde{V}_{3}$ to $\widetilde{I}_{3}$ do disappear because of the intervention on $\widetilde{V}_{3}$ will make the association perturbs.
Experimental Setup
For the FAIR-NN implementation using Gumbel approximation, we also run gradient descent ascent using the Adam optimizer using a learning rate of 1e-3, batch size $64$. The number of iterations is $100k$. In each iteration, one gradient descent update of the neural network parameters in $g$ and the Gumbel logits parameter $w$ is conducted followed by three gradient ascent updates of the neural network parameters in $f^{(0)}$ and $f^{(1)}$. We also use fixed $\gamma=36$. The neural network architectures for all the estimators are the same and are the same as in the simulation of FAIR-NN. The implementation details for all the estimators are:
  • (1)
    Pooled-NN: it simply runs least squares on the full covariate $X$ using all the data.
  • (2)
    FAIR-NN-GB: Our FAIR-NN estimator with Gumbel approximation, its prediction on the test dataset is evaluated by averaging the predictions over $100$ Gumbel samples.
  • (3)
    FAIR-NN-RF: it first selects the variables $x_{j}$ in the fitted model in (2) with ${\mathrm{sig}(w_{j})}>0.9$, i.e., $\widehat{S}=\{j:{\mathrm{sig}(w_{j})}>t\}$, and runs least squares again on $X_{\widehat{S}}$ using all the data.
  • (4)
    Oracle-NN: it runs least squares on $X_{S^{\star}}$ using all the data and neural networks.
  • (5)
    Oracle-Linear: it runs least squares on $X_{S^{\star}}$ using all the data and linear model.
The out-of-sample $R^{2}$ for all the estimators is reported based on the model selection using the validation set that is sampled from the same source as training data with sample size $n^{\prime}=0.6n$. Such a model selection is adopted to prevent the model from over-fitting.
C.4 Details of the Prediction Based on Extracted Features
We generate datasets by combining the bird images in the CUB dataset (Wah et al.,, 2011) and the background images in the Places dataset (Zhou et al.,, 2017) using specific probabilities, which is similar to the waterbird setting in Sagawa et al., (2020) except the spurious correlation ratio. In each environment, there are $50\%$ water birds and $50\%$ land birds. The probabilities of each environment are as follows:
  • (a)
    Environment-1. We place $95\%$ of all water birds against a water background, with the remaining $5\%$ against a land background. We place $90\%$ of all land birds against a land background, with the remaining $10\%$ against a water background. The dataset is denoted by $\mathcal{D}_{1}$, with $50$k images.
  • (b)
    Environment-2. We place $75\%$ of all waterbirds against a water background, with the remaining $25\%$ against a land background. We place $70\%$ of all landbirds against a land background, with the remaining $30\%$ against a water background. The dataset is denoted by $\mathcal{D}_{2}$, with $50$k images.
  • (c)
    Environment-3 (Test Environment). We only place $2\%$ of all waterbirds against a water background, with the remaining $98\%$ against a land background. We place $2\%$ of all landbirds against a land background, with the remaining $98\%$ against a water background. The dataset is denoted by $\mathcal{D}_{3}$, with $30$k images.
  • (d)
    Environment-4 (Oracle Environment). We place $50\%$ of all waterbirds against a water background, with the remaining $50\%$ against a land background. We place $50\%$ of all landbirds against a land background, with the remaining $50\%$ against a water background. The dataset is denoted by $\mathcal{D}_{4}$, with $30$k images.
Class Identification
We apply the CUB dataset Wah et al., (2011), which contains images of birds, along with pixel-level segmentation masks for each bird. When generating the dataset, we classify each bird into waterbird if it belongs to the seabird or waterfowl categories (e.g., albatross, auklet, cormorant, frigatebird, fulmar, gull, jaeger, kittiwake, pelican, puffin, tern, gadwall, grebe, mallard, merganser, guillemot, or Pacific loon) and the land birds if it does not belong to the seabird or waterfowl categories.
Image Generation
When picking bird images from the CUB dataset, we use the provided pixel-level segmentation masks to crop each bird from its original background. Then we decide which environment they should be placed in and select either a water background like ocean and lake or a land background like bamboo forest and broadleaf forest sourced from the Places dataset Zhou et al., (2017). We randomly select $70\%$ of the images in the CUB dataset as a training set and the remaining $30\%$ as a testing set and generate our dataset for training and testing based on the split CUB dataset.
Feature Extraction
Based on the dataset, we use the Pytorch torchvision implementation of the ResNet50 model He et al., (2016) with the pre-trained weights to extract the feature of the images, obtaining a dataset of the feature vector of 2048 dimensions. Then we apply principal components analysis (PCA) to reduce the dimensions of the feature vector to $500$ based on the whole training data $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. We apply the same dimensionality reduction transformation to data in other environments.
Experiment Setup
We run FAIR-Linear with Gumbel approximation on the dataset. Following the standard setting, we apply the logistic loss and Adam optimizer using a learning rate of $21e-$ 2, weight decay of $41e-$ 4, and batch size $4096$ for $10000$ iterations. In each iteration, one gradient descent update of the neural network parameters in $g$ and the Gumbel logits parameter $\omega$ is conducted based on $5$ gradient ascent updates of the neural network parameters in $f$. We fix $\gamma$ as $200$. The implementation details for all the estimators are:
(1) Oracle: it runs logistic regression with $\ell_{1}$ penalty and penalty weight $\alpha=0.001$ on the oracle environment for $1000$ iterations.
(2) Pooled Lasso: it runs logistic regression with $\ell_{1}$ penalty and penalty weight $\alpha=0.001$ on the Environment-1 and Environment-2 for $1000$ iterations.
(3) Lasso on D2: it runs logistic regression with $\ell_{1}$ penalty and penalty weight $\alpha=0.001$ on the Environment-2 for $1000$ iterations.
(4) FAIR-GB: Our FAIR-Linear estimator with Gumbel approximation trained on Environment-1 and Environment-2 for $10000$ iterations.
(5) IRM: it runs Invariant Risk Minimization (IRM) trained on Environment-1 and Environment-2 with $\ell_{2}$ regularizer weight $0.001$ and penalty weight $100$ for $10000$ iterations.
(6) GroupDRO: it runs Group Distributionally Robust Optimization (Group-DRO) on Environment-1 and Environment-2 using ResNet50 and $\gamma=0.1$ for $10000$ iterations.